# Flexible Deep Neural Networks for Partially Linear Survival Data

### v1

@@ Abstract
We propose a flexible deep neural network (DNN) framework for modeling survival data within a partially linear regression structure. The approach preserves interpretability through a parametric linear component for covariates of primary interest, while a nonparametric DNN component captures complex time–covariate interactions among nuisance variables. We refer to the method as FLEXI–Haz, a FLEXIble Hazard model with a partially linear structure. In contrast to existing DNN approaches for partially linear Cox models, FLEXI–Haz does not rely on the proportional hazards assumption. We establish theoretical guarantees: the neural network component attains minimax-optimal convergence rates that depends on composite Hölder classes, and the linear estimator is $\sqrt{n}$–consistent, asymptotically normal, and semiparametrically efficient. Extensive simulations and real-data analyses demonstrate that FLEXI–Haz provides accurate estimation of the linear effect, offering a principled and interpretable alternative to modern methods based on proportional hazards. Code for implementing FLEXI-Haz, as well as scripts for reproducing data analyses and simulations is available at GitHub site https://github.com/AsafBanana/FLEXI-Haz.
keywords: Deep neural networks; Interpretability; Partial linear regression; Survival analysis; Time to event data.
@@ 1 Introduction
Deep neural networks (DNNs) are widely used in modern data analysis because of their ability to capture complex nonlinear patterns, handle high-dimensional inputs, and scale to large datasets. They have achieved strong performance in diverse domains such as image recognition, speech processing, and natural language understanding, often improving on traditional statistical and machine learning methods (lecun2015deep). Continued progress in architectures, optimization, and computational resources has further expanded their range of applications in both scientific and industrial domains.
Deep neural networks for survival analysis.
@@
Survival analysis aims to model the distribution of event times in the presence of censoring, with applications in medicine, public health, epidemiology, engineering, and finance (kalbfleisch2011; klein2006survival). In recent years, DNNs have become increasingly popular in survival analysis thanks to their ability to model complex nonlinear relationships between covariates and event times and to scale efficiently to high-dimensional data. Many existing approaches extend the classical semi-parametric Cox proportional hazards (PH) model (cox1972regression) by replacing the parametric linear risk score with a neural network, while using the negative partial likelihood as the network’s loss function. Early work by faraggi1995neural employed a single hidden-layer network in place of the linear component, and katzman2018deepsurv extended this to deeper architectures (DeepSurv), demonstrating improved concordance in several benchmark datasets. Variants of this approach have been proposed for genomics (yousefi2017predicting; ching2018coxnnet), medical imaging (haarburger2019image; li2019deep), and recurrent or discrete-time formulations (giunchiglia2018rnnsurv; gensheimer2019nnet). While these PH-based DNN methods can yield strong empirical performance, they inherit the PH restriction, which may limit their ability to capture rich time-dependent effects.
Non-PH modeling with deep learning.
To address the limitations of the PH assumption, a growing body of work has focused on DNN-based models that allow covariate effects to vary with time. One strategy augments the network input with time itself, enabling non PH while retaining an unspecified baseline hazard, as in Cox-Time of kvamme2019time. Other approaches abandon the Cox structure entirely in favor of modeling the event time distribution or hazard function directly. In particular, hu2023conditional proposed a framework for estimating the full survival function, using a likelihood function derived from the continuous-time hazard without introducing a baseline hazard function. Their formulation naturally accommodates time-varying covariates through a data-expansion scheme and applies optimization to the log-hazard. By sidestepping the baseline hazard and proportionality constraints, such approaches offer greater modeling flexibility and can better reflect complex temporal dynamics in the hazard. The utility of this approach is illustrated through simulations and real-data analyses, but no asymptotic results are provided.
Semiparametric models and deep learning.
Alongside fully nonparametric formulations, there is considerable interest in semiparametric survival models that combine a parametric component for covariates of primary interest with a flexible, nonparametric component for high-dimensional or nuisance covariates. This structure preserves interpretability and enables formal inference for the parametric component, while leveraging modern machine learning to estimate the nuisance function. Theoretical foundations for this paradigm have been established in the non-survival setting by schmidthieber2020nonparametric, who derived minimax-optimal rates for ReLU-activated DNNs in nonparametric regression, and extended to survival analysis by zhong2022deep in the partially linear Cox model. In the latter, the log-hazard at time $t$ is modeled as $\log h_{0}(t)+\mbox{${\zeta}$}^{\top}{\bf Z}+f({\bf X})$, where $h_{0}(t)$ is an unspecified baseline hazard function, ${\bf Z}\in\mathbb{R}^{p}$ is the vector of covariates of primary interest, ${\bf X}\in\mathbb{R}^{d}$ is the vectors of nuisance covariates, and $\mbox{${\zeta}$}\in\mathbb{R}^{p}$ is the main parameter of interest, estimated at the semiparametric efficiency bound via partial likelihood. The nuisance function $f$ is approximated by a DNN. This approach achieves both optimal convergence for $g$ and $\sqrt{n}$-consistency for $\zeta$} under the PH assumption. The work of zhong2022deep demonstrates how semiparametric theory and deep learning can be integrated to yield interpretable, efficient estimators in survival analysis, but like other Cox-based methods, it remains tied to PH and baseline hazard estimation.
Contributions.
@@
In this work, we develop a semiparametric survival model that extends the partially linear framework of zhong2022deep to the non-PH setting. We model the log-hazard function at time $t$ as $\mbox{${\theta}$}_{o}^{\sf T}{\bf Z}+g_{o}(t,{\bf X})$, where $\mbox{${\theta}$}_{o}\in\mathbb{R}^{p}$ is the parameter of primary interest and $g_{o}$ is an unknown nuisance function approximated by a DNN. This formulation allows arbitrary interactions between the nuisance covariates $\bf X$} and time $t$, naturally accommodates time-dependent effects and time-dependent covariates within $\bf X$}, and does not require specifying or estimating a baseline hazard function.
A central methodological difference from existing semiparametric approaches is that we work with the full likelihood, rather than the partial likelihood commonly used in Cox-type models. Despite removing the PH structure, we establish that the estimator of the parametric component $\mbox{${\theta}$}_{o}$ remains $\sqrt{n}$-consistent, asymptotically normal, and semiparametric efficient. In particular, we show that the theoretical guarantees of zhong2022deep, developed for the partially linear Cox model, continue to hold in the substantially more general non-PH setting, provided the hazard is almost surely bounded away from zero.
While parts of our analysis build on the approximation theory of schmidthieber2020nonparametric and the high-level strategy of zhong2022deep, several key components of our asymptotic theory are fundamentally different. The core technical challenges arise from the fact that our neural network estimator is a function of both $t$ and $\bf X$}, rather than $\bf X$} alone, and the fact that we use the full likelihood rather than a profiling argument. These differences require new empirical process tools, a new characterization of the nuisance tangent space, and a new argument for controlling the bias induced by the DNN approximation.
Finally, when the linear effect $\mbox{${\theta}$}_{o}^{\top}{\bf Z}$ is removed and $\bf Z$} is absorbed into $g_{o}$, our model collapses to the conditional hazard model of hu2023conditional. Our analysis therefore fills an important theoretical gap by providing the first asymptotic guarantees for that model as well.
Code for implementing FLEXI-Haz, as well as scripts for reproducing all data analyses and simulations, is available at GitHub site https://github.com/AsafBanana/FLEXI-Haz.
@@ 2 Model Formulation
Under the above setting, the Cox PH model (cox1972regression) specifies the instantaneous hazard function at time $t>0$ as
@@
where $h_{0}$ is an unspecified baseline hazard function and ${\mbox{${\alpha}$}}_{1}\in\mathbb{R}^{d}$ and ${\mbox{${\alpha}$}}_{2}\in\mathbb{R}^{p}$ are unknown parameters. The model remains widely used due to its interpretability, semiparametric structure, and the availability of efficient estimation procedures. DeepSurv (katzman2018deepsurv) extends the Cox model by replacing the linear risk score $\exp({\mbox{${\alpha}$}}_{1}^{\top}\mathbf{X}+{\mbox{${\alpha}$}}_{2}^{\top}\mathbf{Z})$ with a flexible nonlinear function $f_{1}({{\bf X}},{{\bf Z}})$ estimated by a DNN. Cox-Time (kvamme2019time) further generalized this formulation by replacing $f_{1}$ with a time-depend function, $f_{2}(t,{{\bf X}},{{\bf Z}})$, thereby relaxing the PH assumption and permitting time-covariate interactions. Because $f_{2}$ is time-dependent while the baseline hazard is left unspecified, identifiability becomes nontrivial; kvamme2019time address this challenge through a numerical normalization strategy. More recently, hu2023conditional proposed a baseline-free hazard model of the form
@@
which provides full flexibility for modeling non-proportional, time-varying effects via a DNN. Their estimation procedure is evaluated only through simulations, with no accompanying asymptotic guarantees. Although these DNN-based approaches are highly flexible, they do not provide interpretable estimates for the effect of $\bf Z$}}. To address this limitation, zhong2022deep introduced the deep partially linear Cox model
@@
which augments the Cox framework by separating time component from the covariates and retaining a parametric term $\mbox{${\zeta}$}^{\top}\mathbf{Z}$ for interpretability, while modeling nuisance effects through a DNN. This structure enables valid inference for $\zeta$} while still allowing complex, nonlinear effect of $\bf X$}}.
In this work, we extend the deep partially linear framework of zhong2022deep to the non-PH setting by considering a baseline-free model of the form
$h_{o}(t\mid\mathbf{X},{{\bf Z}})=\exp\{\mbox{${\theta}$}_{o}^{\top}\mathbf{Z}+g_{o}(t,\mathbf{X})\},$ (1)
where $g_{o}(t,\mathbf{X})$ is estimated by a DNN. Our formulation retains the interpretability and inference guarantees for the parametric component ${\mbox{${\theta}$}}_{o}$, but generalizes the nuisance component from $f(\mathbf{X})$ to $g_{o}(t,\mathbf{X})$, thereby allowing fully flexible, time-varying effects and removing the reliance on a baseline hazard function. This extension overcomes the PH restriction inherent in zhong2022deep, resolves the non-identifiability issues that arise when time is included in the DNN, and greatly expands the class of survival models that can be handled within a semiparametric, partially linear framework.
As we demonstrate in Section 6, restricting the nuisance function to depend only on only on $\mathbf{X}$, as in the partially linear Cox model, can induce substantial bias in the estimation of the linear effect whenever the nuisance covariates exhibit time-varying effects. Allowing $g_{o}$ to depend on both $t$ and $\mathbf{X}$ eliminates this source of bias but introduces significant methodological challenges. First, the negative partial likelihood can no longer be used as the DNN loss function. Instead, the full-likelihood is adopted which includes the predictable cumulative hazard that must be numerically approximated at every training step, increasing the computational burden. Second, the presence of this integral markedly complicates the asymptotic analysis: the time dimension introduces additional stochastic terms that must be controlled, requiring new empirical process arguments beyond those used for time-invariant nuisance functions. Our estimation framework in Section 3 addresses these computational difficulties, and the technical developments leading to the asymptotic results are presented in Section 4.3.
@@ 2.1 Composite Hölder Class of $g_{o}$
To formalize the complexity assumptions on the nuisance component $g_{o}$, we follow the compositional function framework of juditsky2009 and schmidthieber2020nonparametric, and use the notation of zhong2022deep. We assume that the nuisance function $g_{o}$ admits a hierarchical, low-dimensional representation of the form
@@
with the convention that $(g\circ f)(x)=g(f(x))$. This composition is associate with a dimension vector ${\bf d}^{*}=(d^{*}_{0},\ldots,d^{*}_{q+1})\in\mathbb{N}_{+}^{q+2}$ and an intrinsic dimension vector $\tilde{\bf d}=(\tilde{d}_{0},\ldots,\tilde{d}_{q})\in\mathbb{N}_{+}^{q+1}$, $\tilde{d}_{\ell}\leq d^{*}_{\ell}$ for each $\ell$. The outermost input is $d^{*}_{0}=r=1+d$, corresponding to observed times and covariates $\mathbf{X}$, and the final output is scalar $d^{*}_{q+1}=1$. For each layer $l$, $l=0,\ldots,q$, the vector-valued mapping $g_{\ell}=(g_{\ell 1},\ldots,g_{\ell\,d^{*}_{\ell+1}})^{\top}$ maps the bounded set $a_{\ell},b_{\ell}]^{\,d^{*}_{\ell}}$ into $a_{\ell+1},b_{\ell+1}]^{\,d^{*}_{\ell+1}}$, where $a_{l},b_{l}\in\mathbb{R}$ are fixed finite constants ensuring each intermediate representation remains bounded. Additionally, each component $g_{\ell j}$ depending on at most $\tilde{d}_{\ell}$ coordinates of its input.
@@
We also assume that each component function satisfies a standard Hölder–smoothness condition. Let $\alpha>0$ and $M>0$. A function is said to have Hölder smoothness index $\alpha$ if all mixed partial derivatives up to order $\lfloor\alpha\rfloor$ exist and are uniformly bounded, and if the derivatives of order exactly $\lfloor\alpha\rfloor$ are $(\alpha-\lfloor\alpha\rfloor)$–Hölder continuous. Here, $\lfloor\alpha\rfloor$ denotes the greatest integer strictly smaller than $\alpha$. The ball of $\alpha$-Hölder functions with radius $M$ consists of all such functions whose derivatives and Hölder constants are bounded by $M$. Formally, for a domain $\mathbb{D}\subset\mathbb{R}^{m}$, the Hölder ball is defined as
$\mathcal{H}^{\alpha}_{m}(\mathbb{D},M)=\Big\{g:\mathbb{D}\to\mathbb{R}:\sum_{\boldsymbol{\beta}:|\boldsymbol{\beta}|<\alpha}\|\partial^{\boldsymbol{\beta}}g\|_{\infty}+\sum_{\boldsymbol{\beta}:|\boldsymbol{\beta}|=\lfloor\alpha\rfloor}\sup_{u,v\in\mathbb{D},u\neq v}\frac{|\partial^{\boldsymbol{\beta}}g(u)-\partial^{\boldsymbol{\beta}}g(v)|}{\|u-v\|^{\alpha-\lfloor\alpha\rfloor}}\leq M\Big\},$
where $\boldsymbol{\beta}^{\sf T}=(\beta_{1},\ldots,\beta_{m})\in\mathbb{N}^{m}$ is a multi-index $\partial^{\boldsymbol{\beta}}=\partial^{\beta_{1}}\cdots\partial^{\beta_{m}}$, $|\boldsymbol{\beta}|=\sum_{j=1}^{m}\beta_{j}$, and $\|g\|_{\infty}$ is the sup–norm of a function over its domain.
@@
$\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)=\Big\{g=g_{q}\circ\cdots\circ g_{0}:g_{\ell j}\in\mathcal{H}^{\alpha_{\ell}}_{\tilde{d}_{\ell}}([a_{\ell},b_{\ell}]^{\tilde{d}_{\ell}},M),\ |a_{\ell}|,|b_{\ell}|\leq M\Big\}.$
This class accomodates a broad range of structured functions, such as additive models, single- and multiple-index models, and hierarchical interactions, through the intrinsic dimensions $\tilde{d}_{\ell}$. Finally, we assume that the true nuisance function $g_{o}$ belongs to the composite class, $g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
In the classical (non-composite) Hölder class $\mathcal{H}^{\alpha}_{d}$, the minimax-optimal rate is $n^{-\alpha/(2\alpha+d)}$ (tsybakov2009). For multilayer composite Hölder classes, schmidthieber2020nonparametric established matching minimax lower and upper bounds (up to polylogarithmic factors), showing that the optimal rate is governed by $\gamma_{n}\;=\;\max_{\ell=0,\ldots,q}\,n^{-\tilde{\alpha}_{\ell}/(2\tilde{\alpha}_{\ell}+\tilde{d}_{\ell})}$, where the effective smoothness indices are $\tilde{\alpha}_{\ell}\;=\;\alpha_{\ell}\prod_{k=\ell+1}^{q}(\alpha_{k}\wedge 1),\ \ a\wedge b=\min\{a,b\}$. Thus, each layer contributes to the overall difficulty of the estimation problem through its intrinsic dimension $\tilde{d}_{\ell}$ and effective smoothness $\tilde{\alpha}_{\ell}$. In Theorem 2 we show that our estimator achieves this optimal composite rate $\gamma_{n}$, up to polylogarithmic factors.

### v2

@@ Abstract
We propose a flexible deep neural network (DNN) framework for modeling survival data within a partially linear regression structure. The approach preserves interpretability through a parametric linear component for covariates of primary interest, while a nonparametric DNN component captures complex time–covariate interactions among nuisance variables. We refer to the method as FLEXI–Haz, a FLEXIble Hazard model with a partially linear structure. In contrast to existing DNN approaches for partially linear Cox models, FLEXI–Haz does not rely on the proportional hazards assumption. We establish theoretical guarantees: the neural network component attains minimax-optimal convergence rates over composite Hölder classes, the linear estimator is $\sqrt{n}$-consistent, asymptotically normal, and semiparametrically efficient, and we develop a cross-fitted one-step estimator of the cumulative hazard and survival function for a new subject, together with pointwise asymptotic confidence intervals. To the best of our knowledge, this is the first frequentist asymptotic pointwise inference result for a survival function in a DNN survival model, with or without a linear component. Simulations and real-data analyses demonstrate the utility of FLEXI–Haz as a principled and interpretable alternative to methods based on proportional hazards.
Time to event data,
keywords:
[class=MSC]
keywords:
[…]
@@ 1 Introduction
Deep neural networks (DNNs) have become a central tool in modern data analysis because of their ability to capture complex nonlinear structure, accommodate high-dimensional inputs, and scale to large data sets. They have achieved strong empirical performance across a wide range of domains, including image recognition, speech processing, and natural language understanding (LeCun et al., 2015). These developments have also stimulated substantial interest in using DNNs for survival analysis.
@@
Survival analysis concerns the distribution of event times in the presence of censoring and arises in many fields, including medicine, public health, epidemiology, engineering, and finance (Kalbfleisch and Prentice, 2011; Klein and Moeschberger, 2006). In recent years, DNN-based survival methods have become increasingly popular because they can model complex nonlinear relationships between covariates and event times while remaining computationally feasible in high dimensions. A large part of this literature extends the Cox proportional hazards (PH) model (Cox, 1972) by replacing the linear risk score with a neural network and training by the negative partial likelihood. Early work by Faraggi and Simon (1995) used a single hidden-layer network, and Katzman et al. (2018) extended this idea to deeper architectures in DeepSurv. Related PH-based DNN methods have been developed for genomics (Yousefi et al., 2017; Ching et al., 2018), medical imaging (Haarburger et al., 2019; Li et al., 2019), recurrent and discrete-time settings (Giunchiglia et al., 2018; Gensheimer and Narasimhan, 2019), and hypothesis testing for covariate effects on survival (Zhong et al., 2026). While these approaches can perform well empirically, they inherit the PH restriction, which may be too rigid when covariate effects vary over time.
To move beyond the PH assumption, several DNN-based methods allow time-varying effects. One approach augments the network input with time while retaining an unspecified baseline hazard, as in Cox-Time (Kvamme et al., 2019). Other approaches use baseline-free formulations based on the continuous-time hazard; see, for example, Hu and Nan (2023); Su et al. (2026). These methods naturally accommodate time-varying effects and avoid the PH constraint, but Hu and Nan (2023) do not provide asymptotic theory, and Su et al. (2026) focus on one- and two-sample hypothesis testing.
At the same time, there is strong interest in semiparametric survival models that retain a low-dimensional interpretable component for covariates of primary interest while modeling high-dimensional or nuisance covariates flexibly. This partly linear structure is attractive because it combines interpretability with statistical efficiency. In the non-survival setting, theoretical foundations for DNN estimation over structured nonparametric classes were developed by Schmidt-Hieber (2020). In survival analysis, Zhong et al. (2022) extended this paradigm to the partially linear Cox model $\log h_{0}(t)+\zeta^{\top}{\bf Z}+f({\bf X})$, where $\bf Z$} is the low-dimensional covariate vector of primary interest, $\bf X$} is a nuisance covariate vector, and $f$ is approximated by a DNN. Their approach yields optimal convergence of the nuisance estimator and $\sqrt{n}$-consistent estimation of $\zeta$, but it remains tied to the PH assumption and to baseline hazard estimation.
@@
In this paper we develop a flexible semiparametric survival model that extends the partially linear framework of Zhong et al. (2022) to a substantially more general non-PH setting. Specifically, we model the log-hazard as $\mbox{${\theta}$}_{o}^{\top}{\bf Z}+g_{o}(t,{\bf X})$, where $\mbox{${\theta}$}_{o}\in\mathbb{R}^{p}$ is the finite-dimensional parameter of interest and $g_{o}(t,{\bf X})$ is an unknown nuisance function approximated by a DNN. This formulation is baseline-free: all time dependence is absorbed into $g_{o}(t,{\bf X})$, rather than decomposed into a baseline hazard and covariate effects. As a result, the model allows arbitrary interactions between time and the nuisance covariates, naturally accommodates time-varying effects, and strictly enlarges the class of partially linear Cox-type models.
This extension is methodologically nontrivial. Once the nuisance component depends on both $t$ and $\bf X$}, the Cox partial likelihood is no longer available as a loss function, and the full likelihood must be used instead. This introduces a predictable cumulative-hazard integral that must be numerically approximated during training and leads to a substantially different asymptotic analysis. In particular, the time-dependent nuisance component requires new empirical-process arguments, a new characterization of the nuisance tangent space, and a new control of the bias induced by DNN approximation. We show, nevertheless, that the estimator of $\mbox{${\theta}$}_{o}$ remains $\sqrt{n}$-consistent, asymptotically normal, and semiparametrically efficient.
A second and particularly important contribution of the paper is inference for the survival curve of a new subject. For an independent draw $({\bf X}_{0},{\bf Z}_{0})$ from the covariate distribution, we construct a cross-fitted one-step estimator of the cumulative hazard $H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$ and the survival function $S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$, together with Wald-type pointwise confidence intervals. To the best of our knowledge, this is the first frequentist asymptotic pointwise inference result for a survival curve in a DNN survival model, with or without a separate linear component. This result is established under an additional assumption that the nuisance covariate vector $\bf X$} has discrete finite support, which yields an explicit closed-form Riesz representer and therefore a tractable one-step correction. From a practical perspective, this restriction is often acceptable because nuisance covariates frequently include categorical variables or grouped continuous features, and because the finite-support assumption is required only for the explicit survival-inference theory rather than for estimation and efficiency of $\mbox{${\theta}$}_{o}$. The broader survival-inference problem can also be studied under general covariate distributions through a high-level semiparametric characterization of the efficient influence function. Since that formulation is more technical and does not yield a comparably explicit estimator without additional approximation arguments, we present it in the appendix.
Our framework is also relevant beyond the partially linear setting. If the linear component is absorbed into the nuisance function, then the one-step construction and efficient influence function derived here remain valid, with the parametric term dropping out. In this sense, the paper also contributes asymptotic survival-inference tools for baseline-free DNN hazard models more broadly. Compared with Su et al. (2026), who derive functional asymptotic normality for smooth hazard-function functionals and use it for testing, we study a different and sharper target: semiparametric efficiency for an interpretable linear effect and explicit Wald-type pointwise inference for the survival curve of a new subject.
The remainder of the paper is organized as follows. Sections 2 and 3 introduce the model, the DNN class, and the FLEXI–Haz estimation and survival-inference procedures. Sections 4 and 5 develop the asymptotic theory for estimation and pointwise survival inference. Simulation studies and real-data analyses are presented later. Code for FLEXI–Haz and for reproducing the simulations and data analyses is available at https://github.com/AsafBanana/FLEXI--Haz.
@@ 2 Model Formulation
Under the above setting, the Cox PH model (Cox, 1972) specifies the instantaneous hazard function at time $t>0$ as
@@
where $h_{0}$ is an unspecified baseline hazard function and ${\mbox{${\alpha}$}}_{1}\in\mathbb{R}^{d}$ and ${\mbox{${\alpha}$}}_{2}\in\mathbb{R}^{p}$ are unknown parameters. The model remains widely used due to its interpretability, semiparametric structure, and the availability of efficient estimation procedures. DeepSurv (Katzman et al., 2018) extends the Cox model by replacing the linear risk score $\exp({\mbox{${\alpha}$}}_{1}^{\top}\mathbf{X}+{\mbox{${\alpha}$}}_{2}^{\top}\mathbf{Z})$ with a flexible nonlinear function $f_{1}({{\bf X}},{{\bf Z}})$ estimated by a DNN. Cox-Time (Kvamme et al., 2019) further generalized this formulation by replacing $f_{1}$ with a time-dependent function, $f_{2}(t,{{\bf X}},{{\bf Z}})$, thereby relaxing the PH assumption and permitting time-covariate interactions. Because $f_{2}$ is time-dependent while the baseline hazard is left unspecified, identifiability becomes nontrivial; Kvamme et al. (2019) address this challenge through a numerical normalization strategy. More recently, Hu and Nan (2023) proposed a baseline-free hazard model of the form
@@
which provides full flexibility for modeling non-proportional, time-varying effects via a DNN. Their estimation procedure is evaluated only through simulations, with no accompanying asymptotic guarantees. Although these DNN-based approaches are highly flexible, they do not provide interpretable estimates for the effect of $\bf Z$}}. To address this limitation, Zhong et al. (2022) introduced the deep partially linear Cox model
@@
which augments the Cox framework by separating the time component from the covariates and retaining a parametric term $\mbox{${\zeta}$}^{\top}\mathbf{Z}$ for interpretability, while modeling nuisance effects through a DNN. This structure enables valid inference for $\zeta$} while still allowing complex, nonlinear effect of $\bf X$}}.
In this work, we extend the deep partially linear framework of Zhong et al. (2022) to the non-PH setting by considering a baseline-free model of the form
$h_{o}(t\mid\mathbf{X},{{\bf Z}})=\exp\{\mbox{${\theta}$}_{o}^{\top}\mathbf{Z}+g_{o}(t,\mathbf{X})\},$ (2.1)
where $g_{o}(t,\mathbf{X})$ is estimated by a DNN. Model (2.1) is baseline-free in the sense that the entire time dependence of the hazard is absorbed into the nuisance function $g_{o}(t,{\bf X})$, rather than being decomposed into a baseline hazard and covariate effects. This is not merely a reparameterization of a Cox-type model with time-varying coefficients: allowing $g_{o}(t,{\bf X})$ to be an unrestricted function of $(t,{\bf X})$ strictly enlarges the model class relative to representations of the form $h_{0}(t)\exp\{\zeta^{\top}{\bf Z}+f(X)\}$. This extension overcomes the PH restriction inherent in Zhong et al. (2022), resolves the non-identifiability issues that arise when time is included in the DNN, and greatly expands the class of survival models that can be handled within a semiparametric, partially linear framework. Identifiability of $\mbox{${\theta}$}_{o}$ in this setting follows from the partially linear structure together with Assumption A5 (see Section 4.2). These conditions ensure that no nontrivial linear function of $\bf Z$} can be represented almost surely as a function of $(t,{\bf X})$, and therefore prevent $g_{o}(t,{\bf X})$ from absorbing components of the linear effect. This is formulated in Lemma 3 (Section 4.3).
As we demonstrate in Section 7, restricting the nuisance function to depend only on $\mathbf{X}$, as in the partially linear Cox model, can induce substantial bias in the estimation of the linear effect whenever the nuisance covariates exhibit time-varying effects. Allowing $g_{o}$ to depend on both $t$ and $\mathbf{X}$ eliminates this source of bias but introduces significant methodological challenges. First, the negative partial likelihood can no longer be used as the DNN loss function. Instead, the full likelihood is adopted, which includes the predictable cumulative hazard that must be numerically approximated at every training step, increasing the computational burden. Second, the presence of this integral markedly complicates the asymptotic analysis: the time dimension introduces additional stochastic terms that must be controlled, requiring new empirical process arguments beyond those used for time-invariant nuisance functions. Our estimation framework in Section 3 addresses these computational difficulties, and the technical developments leading to the asymptotic results are presented in Sections 4 and 5.
@@ 2.1 Composite Hölder Class of $g_{o}$
To formalize the complexity assumptions on the nuisance component $g_{o}$, we follow the compositional function framework of Juditsky et al. (2009) and Schmidt-Hieber (2020), and use the notation of Zhong et al. (2022). We assume that the nuisance function $g_{o}$ admits a hierarchical, low-dimensional representation of the form
@@
with the convention that $(g\circ f)(x)=g(f(x))$. This composition is associated with a dimension vector ${\bf d}^{*}=(d^{*}_{0},\ldots,d^{*}_{q+1})\in\mathbb{N}_{+}^{q+2}$ and an intrinsic dimension vector $\tilde{\bf d}=(\tilde{d}_{0},\ldots,\tilde{d}_{q})\in\mathbb{N}_{+}^{q+1}$, $\tilde{d}_{\ell}\leq d^{*}_{\ell}$ for each $\ell$. The outermost input is $d^{*}_{0}=r=1+d$, corresponding to observed times and covariates $\mathbf{X}$, and the final output is scalar $d^{*}_{q+1}=1$. For each layer $l$, $l=0,\ldots,q$, the vector-valued mapping $g_{\ell}=(g_{\ell 1},\ldots,g_{\ell\,d^{*}_{\ell+1}})^{\top}$ maps the bounded set $a_{\ell},b_{\ell}]^{\,d^{*}_{\ell}}$ into $a_{\ell+1},b_{\ell+1}]^{\,d^{*}_{\ell+1}}$, where $a_{l},b_{l}\in\mathbb{R}$ are fixed finite constants ensuring each intermediate representation remains bounded. Additionally, each component $g_{\ell j}$ depends on at most $\tilde{d}_{\ell}$ coordinates of its input.
@@
We also assume that each component function satisfies a standard Hölder–smoothness condition. Let $\alpha>0$ and $M>0$. A function is said to have Hölder smoothness index $\alpha$ if all mixed partial derivatives up to order $\lfloor\alpha\rfloor$ exist and are uniformly bounded, and if the derivatives of order exactly $\lfloor\alpha\rfloor$ are $(\alpha-\lfloor\alpha\rfloor)$–Hölder continuous. Here, $\lfloor\alpha\rfloor$ denotes the greatest integer strictly smaller than $\alpha$. The ball of $\alpha$-Hölder functions with radius $M$ consists of all such functions whose derivatives and Hölder constants are bounded by $M$. Formally, for a domain $\mathbb{D}\subset\mathbb{R}^{r}$, the Hölder ball is defined as
$\mathcal{H}^{\alpha}_{r}(\mathbb{D},M)=\Big\{g:\mathbb{D}\to\mathbb{R}:\sum_{\boldsymbol{\beta}:|\boldsymbol{\beta}|<\alpha}\|\partial^{\boldsymbol{\beta}}g\|_{\infty}+\sum_{\boldsymbol{\beta}:|\boldsymbol{\beta}|=\lfloor\alpha\rfloor}\sup_{u,v\in\mathbb{D},u\neq v}\frac{|\partial^{\boldsymbol{\beta}}g(u)-\partial^{\boldsymbol{\beta}}g(v)|}{\|u-v\|^{\alpha-\lfloor\alpha\rfloor}}\leq M\Big\},$
where $\boldsymbol{\beta}^{\sf T}=(\beta_{1},\ldots,\beta_{r})\in\mathbb{N}^{r}$ is a multi-index $\partial^{\boldsymbol{\beta}}=\partial^{\beta_{1}}\cdots\partial^{\beta_{r}}$, $|\boldsymbol{\beta}|=\sum_{j=1}^{r}\beta_{j}$, and $\|g\|_{\infty}$ is the sup–norm of a function over its domain.
@@
$\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)=\Big\{g=g_{q}\circ\cdots\circ g_{0}:g_{\ell j}\in\mathcal{H}^{\alpha_{\ell}}_{\tilde{d}_{\ell}}([a_{\ell},b_{\ell}]^{\tilde{d}_{\ell}},M),\ |a_{\ell}|,|b_{\ell}|\leq M\Big\}.$
This class accommodates a broad range of structured functions, such as additive models, single- and multiple-index models, and hierarchical interactions, through the intrinsic dimensions $\tilde{d}_{\ell}$. Finally, we assume that the true nuisance function $g_{o}$ belongs to the composite class, $g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$.
In the classical (non-composite) Hölder class $\mathcal{H}^{\alpha}_{d}$, the minimax-optimal rate is $n^{-\alpha/(2\alpha+d)}$ (Tsybakov, 2009). For multilayer composite Hölder classes, Schmidt-Hieber (2020) established matching minimax lower and upper bounds (up to polylogarithmic factors), showing that the optimal rate is governed by $\gamma_{n}\;=\;\max_{\ell=0,\ldots,q}\,n^{-\tilde{\alpha}_{\ell}/(2\tilde{\alpha}_{\ell}+\tilde{d}_{\ell})}$, where the effective smoothness indices are $\tilde{\alpha}_{\ell}\;=\;\alpha_{\ell}\prod_{k=\ell+1}^{q}(\alpha_{k}\wedge 1),\ \ a\wedge b=\min\{a,b\}$. Thus, each layer contributes to the overall difficulty of the estimation problem through its intrinsic dimension $\tilde{d}_{\ell}$ and effective smoothness $\tilde{\alpha}_{\ell}$. Taken together, Theorems 1 and 2 show that $\gamma_{n}$ is minimax-optimal up to polylogarithmic factors: Theorem 1 gives the upper bound for our estimator, and Theorem 2 gives the lower bound.